# Computing an Aircraft’s Gliding Range and Minimal Return Altitude in Presence of Obstacles and Wind

### v1

@@ Abstract
In the event of a total loss of thrust a pilot must identify a reachable landing site and subsequently execute a forced landing. To do so this, they must estimate which region on the ground can be reached safely in gliding flight. We call this the gliding reachable region (GRR). To compute the GRR, we employ an optimal control formulation aiming to reach a point in space while minimizing altitude loss. A simplified model of the aircraft’s dynamics is used, where the effect of turns is neglected. The resulting equations are discretized on a grid and solved numerically. Our algorithm for computing the GRR is fast enough to run in real time during flight, it accounts for ground obstacles and wind, and for each point in the GRR it outputs the path to reach it with minimal loss of altitude.
A related problem is estimating the minimal altitude an aircraft needs to glide to a given airfield in the presence of obstacles. This information enables pilots to plan routes that always have an airport within gliding distance. We formalize this problem using an optimal control formulation based on the same aircraft dynamics model. The resulting equations are solved with a second algorithm that outputs the minimal re-entry altitude and the paths to reach the airfield from any position while avoiding obstacles.
@@ I.1 Motivation
The gliding reachable region (GRR) is the region on the ground that can be safely reached in gliding flight from a certain position and altitude. Knowing the GRR allows pilots to quickly assess which landing sites are within reach in the event of an engine failure. To aid in this task several avionics companies such as Garmin, Foreflight and LXNAV have developed respectively the Glide Range Ring [3], Glide Advisor [4] and glide range area [5] aids. These algorithms all display the GRR on the navigation map in real time, allowing the pilot to quickly determine the reachability of a landing site. Aside from emergencies, estimating the reachable region is also of great importance for glider pilots. In this work we consider two problems:
@@
We already commented on the relevance of the first problem, we now comment on why the second one is also interesting. Suppose we want to know if we can reach a certain airport in gliding flight starting from position $\bm{x}\in\operatorname{\mathbb{R}}^{2}$ and altitude $z$. One option is to compute the GRR and check if the airport falls within it. However finding the GRR can be computationally expensive. Suppose instead we have access to a database that for each airfield $q$ and position $\bm{y}\in\operatorname{\mathbb{R}}^{2}$ stores $A_{q}(\bm{y})$, the minimal altitude to return to the airfield from $\bm{y}$. Then to establish if airfield $q$ is reachable it’s sufficient to check if $z\geq A_{q}(\bm{x})$. In principle the contour lines of $A_{q}(\cdot)$ can be printed onto a paper map, allowing the pilot to know if they’re within gliding range of the airfield without resorting to electronic devices. In addition, knowing the minimal return altitudes in advance can help pilots to plan flights and maintain an altitude sufficient to always reach a landable location.
GRRP and MRAP appear to be two quite different problems. We decide to study them in the same article because we employ similar mathematical formalism and algorithms to solve them.
@@ I.2 Our contribution
We develop the algorithms Glikonal-G and Glikonal-M to solve respectively GRRP and MRAP. The algorithms are respectively based on a variant of the Ordered Upwind Method (OUM) [1] and Fast Marching Method [2] (FMM). We claim Glikonal-G is fast enough to run on an on-board computer in real time, taking less than $0.2$ seconds to run on a $100\times 100$ grid. Also given a point within the GRR, Glikonal-G can compute the optimal gliding path to reach it. Similarly Glikonal-M can produce a feasible (but not necessarily optimal, in the sense of minimal altitude loss) path that leads to the airfield.
First we model the aircraft’s dynamics using a 3DOF model, where the airspeed vector is the control parameter. We assume in particular that the aircraft is able to take arbitrarily sharp turns. Also the sink rate of the aircraft is assumed to depend exclusively on the airspeed, thus we do not take into account the fact that an airplane experiences a more negative sink rate when taking turns. In Appendix D we comment on the validity of this approximation and on methods to overcome it.
@@ I.3 Related literature
The simplest method to compute the GRR in absence of obstacles is to assume that the aircraft will follow a path composed by an initial turn towards the desired heading and then a straight path at the best glide speed. This was explored in [6, 7] where the turn with the least altitude loss is computed, and in [8] where the effect of wind is also considered (although without optimizing the turn).
In other works the task of finding a gliding path between two points, where the initial and final headings are also given, is considered. In [9, 10, 8, 11, 12, 13] paths are allowed to be more complex, by concatenating basic maneuvers (e.g., constant bank turns, steady straight flight, accelerated straight flight) in a sequence. In none of these works however obstacles are considered. Most of these studies rely on a generalization of Dubins paths [14], i.e., paths whose maximal curvature is bounded.
@@
In other studies, the optimal control formulation of the problem is also considered. In particular [16, 17] use HJB equations to conduct the reachability analysis. However [17] considers only the longitudinal movements of the aircraft (i.e., no turns are allowed). Another line of work proposed to deal with obstacles by means of a visibility graph [18, 19]. The obstacles are modeled as polygons: in this case the shortest path between two points on the plane, while avoiding the polygons, will touch the polygons exclusively on the vertices. This remark is exploited to build piecewise linear paths that avoid obstacles. The work closest to ours is [20], in which a similar aircraft dynamics to ours is used and obstacles of arbitrary shape are considered. The algorithm in [20] relies on a modified visibility graph to compute the paths with the minimal altitude loss. Glikonal-G does fundamentally the same thing, while being more efficient and accounting for wind. Other approaches to the problem of finding gliding paths between two points rely on genetic algorithms [21], Pseudo-Spectral-Methods [22], dedicated optimization schemes [23].
Except for [6, 8, 7] all other algorithms allow to compute a gliding path between an initial and a single final point. Therefore to compute the GRR one would need to run the algorithm multiple times changing the final point each time, which is unpractical. In contrast Glikonal-G computes all the trajectories to reach points in the GRR in a single run.
@@
Finally let us comment on the origin of the algorithms we use to solve the problem. Sethian first developed the FMM to solve the eikonal equation in [2]. The eikonal equation describes the motion of a propagating front in a medium. As we will see, the anisotropic version of the eikonal equation, belonging to the class of static HJB equations [25, 26], is linked to the optimal dynamics of the aircraft in presence of wind. A set of algorithmic schemes called Ordered Upwind Methods (OUM) able to solve this class of equations has been developed by Sethian and Vladimirsky in [27, 1]. Glikonal-G is a modification of the OUM, instead Glikonal-M is a modified FMM.
@@ II.1 Aircraft dynamics
In this section we introduce the dynamical model for the glider and define the optimal control problem. Throughout the paper we assume all reference systems are East-North-up oriented. Let $\bm{r}=(\bm{x},z)$ indicate the position of the glider in 3D space, where $\bm{x}=(x_{1},x_{2})$ represents the horizontal position, and $z$ is the altitude above mean sea level. Consider also a wind vector field $\bm{W}:\operatorname{\mathbb{R}}^{3}\to\operatorname{\mathbb{R}}^{2}$. $\bm{W}(\bm{r})$ is a two component vector indicating the wind (i.e., the horizontal component of the air velocity with respect to the ground) at position $\bm{r}$. We model the dynamics of the glider in the following way:
@@
Where $\tilde{\bm{a}}(t)$ is the horizontal component of the airspeed at time $t$, and $s:\operatorname{\mathbb{R}}_{+}\to(-\infty,0)$ is the sink rate as a function of the horizontal airspeed.  [^1] We can also model the vertical movements of the air with a function $s_{\text{air}}:\operatorname{\mathbb{R}}^{3}\to\operatorname{\mathbb{R}}$. $s_{\text{air}}(\bm{r})$ is the vertical velocity of the air mass at position $\bm{r}$. For technical reasons the overall sink rate must be negative, otherwise the glider can climb up by staying put on a particular point. To take vertical air movements into account, just replace $s\to s+s_{\text{air}}$ in the rest of the derivation. In straight, steady state gliding flight, to each horizontal airspeed corresponds a vertical airspeed (or sink rate). The functional form of $s$ is given by the polar curve of the aircraft.
@@
The control parameter is $\tilde{\bm{a}}:\operatorname{\mathbb{R}}_{+}\mapsto\operatorname{\mathbb{R}}^{2}$. The only assumption we make about $\tilde{\bm{a}}$ is that it is a measurable function. This implies that we allow the glider’s speed to change discontinuously both in direction and magnitude. $\tilde{\bm{a}}$ is subject to the constraint $V_{\text{S}}<\left\lVert\bm{V}_{\text{air}}\right\rVert<V_{\text{NE}}$, where $V_{\text{S}}$ and $V_{\text{NE}}$ are respectively the stall and never exceed speeds of the aircraft. Let us make some additional remarks about the limitations of our dynamical model. Since the sink rate depends exclusively on the magnitude of the speed and not on the acceleration, the model neglects the effect of turns. In reality the sink rate is higher (more negative) during a turn. Also arbitrarily sharp turns are allowed under this model.
Define the ground speed to be $\bm{a}(t)=\tilde{\bm{a}}(t)+\bm{W}(\bm{r}(t))$. For mathematical convenience we suppose that our control parameter is directly the ground speed $\bm{a}$. Reformulating the equations of motion in terms of $\bm{a}$ we obtain
@@ II.2 Gliding Reachable Region Problem
We now aim to write an optimal control problem equivalent to (3) but where time is replaced with the lost altitude. First remark that since $s<0$ the altitude is a decreasing function of time. Defining the lost altitude $h(t)=z_{0}-z(t)$ we can we can reparametrize the dynamics using $h$ in place of $t$. With a change of variable we obtain
@@
In the second to last step we renamed $\bm{a}(t(h))\to\bm{a}(h)$ and $\bm{W}(\bm{r}(t(h)))\to\bm{W}(\bm{x}(h),z_{0}-h)$. In the last step we introduced function $\tilde{g}(\bm{x},z,\bm{a})\coloneqq\left\lVert\bm{a}\right\rVert/|s(\left\lVert\bm{a}-\bm{W}(\bm{x},z)\right\rVert)|$ representing the glide ratio of the aircraft: it tells us how many meters the aircraft moves horizontally for each meter of lost altitude. Notice $\tilde{g}$ depends on the position of the aircraft and on its ground speed. When the wind is null, $\tilde{g}$ depends only on $\left\lVert\bm{a}\right\rVert$. As a result of this change of variable, we obtained a new formulation of the optimal control problem, where we got rid of the variable $z$. The reparametrized problem is
@@
where $\bm{r}(h)=(\bm{x}(h),z_{0}-h)$. At this point we can derive the HJB equation corresponding to this problem. Let $\bm{x}:[0,h_{f}]\mapsto\operatorname{\mathbb{R}}^{2}$ be the trajectory that minimizes the lost altitude when starting from point $\bm{x}(0)=x_{0},\,z(0)=z_{0}$ and arriving in $\bm{x}(h_{f})=\bm{y},z(h_{f})=z_{0}-h_{f}$. Then we can write
@@
At this point we can further simplify the equation by first changing the minimum into a maximum using $\min_{x}f(x)=-\max_{x}-f(x)$, and then carrying out the optimization with respect to $\left\lVert\bm{a}\right\rVert$, and be left therefore with optimizing only with respect to the direction $\bm{\hat{a}}\coloneqq\bm{a}/\left\lVert\bm{a}\right\rVert$. In formulas
@@
Equation (13) is a first order nonlinear PDE belonging to the well known class of static Hamilton-Jacobi PDEs [2, 27, 1]. In general it does not admit a closed form solution, however very efficient algorithms exist to solve it approximately. Before moving on let us briefly comment on $g$. In our derivation we supposed that the aircraft flies at the speed that yields the optimal glide ratio, implying that the pilot should adapt the magnitude of the airspeed to the local wind. Alternatively, to obtain more conservative estimates one can assume instead that the aircraft flies at a constant horizontal airspeed $v$. In this case $g(\bm{x},z,\bm{\hat{a}})=\left\lVert\tilde{\bm{a}}+\bm{W}(\bm{x},z)\right\rVert/s(v)$, where $\tilde{\bm{a}}$ solves $\bm{\hat{a}}=\frac{\bm{W}(\bm{x},z)+\tilde{\bm{a}}}{\left\lVert\bm{W}(\bm{x},z)+\tilde{\bm{a}}\right\rVert}$, with $\left\lVert\tilde{\bm{a}}\right\rVert=v$.
@@ II.3 Gliding as front propagation
which is known as the eikonal equation. The initial condition is $U(\bm{x}_{0})=z_{0}$. The eikonal equation arises when describing the arrival times of a front in each point in space. Suppose we have a sheet of paper and at time $t=0$ we set fire to it in at a point $\bm{x}_{0}\in\operatorname{\mathbb{R}}^{2}$. Suppose further that the fire front advances through the paper with speed $g$. Then letting $U(\bm{y})$ be the time at which the fire front reaches $\bm{y}$, we have that $U$ satisfies (16). We can also include obstacles in this analogy. Suppose some portion of the paper is wet. In this case, the fire front stops at the border of the wet region. This is similar to what happens when the glider’s altitude falls below $h_{\text{min}}$. This analogy shows how we can see the descending glider as an expanding front that passes around the obstacles until it collides with them and stops.

### v2

@@ Abstract
In the event of a total loss of thrust, a pilot must identify a reachable landing site and subsequently execute a forced landing. To do so, they must estimate which region on the ground can be reached safely in gliding flight. We call this the gliding reachable region (GRR). To compute the GRR, we employ an optimal control formulation aiming to reach a point in space while minimizing altitude loss. A simplified model of the aircraft’s dynamics is used, where the effect of turns is neglected. The resulting equations are discretized on a grid and solved numerically. Our algorithm for computing the GRR is fast enough to run in real time during flight, it accounts for ground obstacles and wind, and for each point in the GRR it outputs the path to reach it with minimal loss of altitude.
A related problem is estimating the minimal altitude an aircraft needs in order to glide to a given airfield in the presence of obstacles. This information enables pilots to plan routes that always have an airport within gliding distance. We formalize this problem using an optimal control formulation based on the same aircraft dynamics model. The resulting equations are solved with a second algorithm that outputs the minimal re-entry altitude and the paths to reach the airfield from any position while avoiding obstacles.
@@ I.1 Motivation
The gliding reachable region (GRR) is the region on the ground that can be safely reached in gliding flight from a certain position and altitude. Knowing the GRR allows pilots to quickly assess which landing sites are within reach in the event of an engine failure. To aid in this task several avionics companies such as Garmin, Foreflight, and LXNAV have developed respectively the Glide Range Ring [3], Glide Advisor [4], and glide range area [5] aids. These algorithms all display the GRR on the navigation map in real time, allowing the pilot to quickly determine the reachability of a landing site. Aside from emergencies, estimating the reachable region is also of great importance for glider pilots. In this work, we consider two problems:
@@
We already commented on the relevance of the first problem, we now comment on why the second one is also interesting. Suppose we want to know if we can reach a certain airport in gliding flight starting from position $\bm{x}\in\operatorname{\mathbb{R}}^{2}$ and altitude $z$. One option is to compute the GRR and check if the airport falls within it. However, finding the GRR can be computationally expensive. Suppose instead we have access to a database that for each airfield $q$ and position $\bm{y}\in\operatorname{\mathbb{R}}^{2}$ stores $A_{q}(\bm{y})$, the minimal altitude to return to the airfield from $\bm{y}$. Then to establish if airfield $q$ is reachable it’s sufficient to check if $z\geq A_{q}(\bm{x})$. In principle the contour lines of $A_{q}(\cdot)$ can be printed onto a paper map, allowing the pilot to know if they’re within gliding range of the airfield without resorting to electronic devices. In addition, knowing the minimal return altitudes in advance can help pilots to plan flights and maintain an altitude sufficient to always reach a landable location.
GRRP and MRAP appear to be two quite different problems. We decide to study them in the same article because we employ similar mathematical formalisms and algorithms to solve them.
@@ I.2 Our contribution
We develop the algorithms Glikonal-G and Glikonal-M to solve respectively GRRP and MRAP. The algorithms are respectively based on a variant of the Ordered Upwind Method (OUM) [1] and Fast Marching Method [2] (FMM). We claim Glikonal-G is fast enough to run on an on-board computer in real time, taking less than $0.2$ seconds to run on a $100\times 100$ grid. Also given a point within the GRR, Glikonal-G can compute the optimal gliding path to reach it. Similarly, Glikonal-M can produce a feasible (but not necessarily optimal, in the sense of minimal altitude loss) path that leads to the airfield.
First, we model the aircraft’s dynamics using a 3DOF model, where the airspeed vector is the control parameter. We assume in particular that the aircraft is able to take arbitrarily sharp turns. Also, the sink rate of the aircraft is assumed to depend exclusively on the airspeed, thus we do not take into account the fact that an airplane experiences a more negative sink rate when taking turns. In Appendix D we comment on the validity of this approximation and on methods to overcome it.
@@ I.3 Related literature
The simplest method to compute the GRR in the absence of obstacles is to assume that the aircraft will follow a path composed by an initial turn towards the desired heading and then a straight path at the best glide speed. This was explored in [6, 7] where the turn with the least altitude loss is computed, and in [8] where the effect of wind is also considered (although without optimizing the turn).
In other works the task of finding a gliding path between two points, where the initial and final headings are also given, is considered. In [9, 10, 8, 11, 12, 13] paths are allowed to be more complex, by concatenating basic maneuvers (e.g., constant bank turns, steady straight flight, accelerated straight flight) in a sequence. In none of these works, however, obstacles are considered. Most of these studies rely on a generalization of Dubins paths [14], i.e., paths whose maximal curvature is bounded.
@@
In other studies, the optimal control formulation of the problem is also considered. In particular [16, 17] use HJB equations to conduct the reachability analysis. However [17] considers only the longitudinal movements of the aircraft (i.e., no turns are allowed). Another line of work proposed to deal with obstacles by means of a visibility graph [18, 19]. The obstacles are modeled as polygons: in this case, the shortest path between two points on the plane, while avoiding the polygons, will touch the polygons exclusively on the vertices. This remark is exploited to build piecewise linear paths that avoid obstacles. The work closest to ours is [20], in which a similar aircraft dynamics to ours is used and obstacles of arbitrary shape are considered. The algorithm in [20] relies on a modified visibility graph to compute the paths with the minimal altitude loss. Glikonal-G does fundamentally the same thing, while being more efficient and accounting for wind. Other approaches to the problem of finding gliding paths between two points rely on genetic algorithms [21], Pseudo-Spectral-Methods [22], dedicated optimization schemes [23].
Except for [6, 8, 7] all other algorithms allow to compute a gliding path between an initial and a single final point. Therefore to compute the GRR one would need to run the algorithm multiple times changing the final point each time, which is unpractical. In contrast, Glikonal-G computes all the trajectories to reach points in the GRR in a single run.
@@
Finally, let us comment on the origin of the algorithms we use to solve the problem. Sethian first developed the FMM to solve the eikonal equation in [2]. The eikonal equation describes the motion of a propagating front in a medium. As we will see, the anisotropic version of the eikonal equation, belonging to the class of static HJB equations [25, 26], is linked to the optimal dynamics of the aircraft in the presence of wind. A set of algorithmic schemes called Ordered Upwind Methods (OUM) able to solve this class of equations has been developed by Sethian and Vladimirsky in [27, 1]. Glikonal-G is a modification of the OUM, instead, Glikonal-M is a modified FMM.
@@ II.1 Aircraft dynamics
In this section, we introduce the dynamical model for the glider and define the optimal control problem. Throughout the paper, we assume all reference systems are East-North-up oriented. Let $\bm{r}=(\bm{x},z)$ indicate the position of the glider in 3D space, where $\bm{x}=(x_{1},x_{2})$ represents the horizontal position, and $z$ is the altitude above mean sea level. Consider also a wind vector field $\bm{W}:\operatorname{\mathbb{R}}^{3}\to\operatorname{\mathbb{R}}^{2}$. $\bm{W}(\bm{r})$ is a two-component vector indicating the wind (i.e., the horizontal component of the air velocity with respect to the ground) at position $\bm{r}$. We model the dynamics of the glider in the following way:
@@
Where $\tilde{\bm{a}}(t)$ is the horizontal component of the airspeed at time $t$, and $s:\operatorname{\mathbb{R}}_{+}\to(-\infty,0)$ is the sink rate as a function of the horizontal airspeed.  [^1] We can also model the vertical movements of the air with a function $s_{\text{air}}:\operatorname{\mathbb{R}}^{3}\to\operatorname{\mathbb{R}}$. $s_{\text{air}}(\bm{r})$ is the vertical velocity of the air mass at position $\bm{r}$. For technical reasons the overall sink rate must be negative, otherwise the glider can climb up by staying put on a particular point. To take vertical air movements into account, just replace $s\to s+s_{\text{air}}$ in the rest of the derivation. In straight, steady-state gliding flight, each horizontal airspeed corresponds to a vertical airspeed (or sink rate). The functional form of $s$ is given by the polar curve of the aircraft.
@@
The control parameter is $\tilde{\bm{a}}:\operatorname{\mathbb{R}}_{+}\mapsto\operatorname{\mathbb{R}}^{2}$. The only assumption we make about $\tilde{\bm{a}}$ is that it is a measurable function. This implies that we allow the glider’s speed to change discontinuously both in direction and magnitude. $\tilde{\bm{a}}$ is subject to the constraint $V_{\text{S}}<\left\lVert\bm{V}_{\text{air}}\right\rVert<V_{\text{NE}}$, where $V_{\text{S}}$ and $V_{\text{NE}}$ are respectively the stall and never exceed speeds of the aircraft. Let us make some additional remarks about the limitations of our dynamical model. Since the sink rate depends exclusively on the magnitude of the speed and not on the acceleration, the model neglects the effect of turns. In reality the sink rate is higher (more negative) during a turn. Also, arbitrarily sharp turns are allowed under this model.
Define the ground speed to be $\bm{a}(t)=\tilde{\bm{a}}(t)+\bm{W}(\bm{r}(t))$. For mathematical convenience, we suppose that our control parameter is directly the ground speed $\bm{a}$. Reformulating the equations of motion in terms of $\bm{a}$ we obtain
@@ II.2 Gliding Reachable Region Problem
We now aim to write an optimal control problem equivalent to (3) but where time is replaced with the lost altitude. First remark that since $s<0$ the altitude is a decreasing function of time. Defining the lost altitude $h(t)=z_{0}-z(t)$ we can we can reparametrize the dynamics using $h$ in place of $t$. With a change of variable, we obtain
@@
In the second to last step we renamed $\bm{a}(t(h))\to\bm{a}(h)$ and $\bm{W}(\bm{r}(t(h)))\to\bm{W}(\bm{x}(h),z_{0}-h)$. In the last step, we introduced function $\tilde{g}(\bm{x},z,\bm{a})\coloneqq\left\lVert\bm{a}\right\rVert/|s(\left\lVert\bm{a}-\bm{W}(\bm{x},z)\right\rVert)|$ representing the glide ratio of the aircraft: it tells us how many meters the aircraft moves horizontally for each meter of lost altitude. Notice $\tilde{g}$ depends on the position of the aircraft and on its ground speed. When the wind is null, $\tilde{g}$ depends only on $\left\lVert\bm{a}\right\rVert$. As a result of this change of variable, we obtained a new formulation of the optimal control problem, where we got rid of the variable $z$. The reparametrized problem is
@@
where $\bm{r}(h)=(\bm{x}(h),z_{0}-h)$. At this point, we can derive the HJB equation corresponding to this problem. Let $\bm{x}:[0,h_{f}]\mapsto\operatorname{\mathbb{R}}^{2}$ be the trajectory that minimizes the lost altitude when starting from point $\bm{x}(0)=x_{0},\,z(0)=z_{0}$ and arriving in $\bm{x}(h_{f})=\bm{y},z(h_{f})=z_{0}-h_{f}$. Then we can write
@@
At this point, we can further simplify the equation by first changing the minimum into a maximum using $\min_{x}f(x)=-\max_{x}-f(x)$, and then carrying out the optimization with respect to $\left\lVert\bm{a}\right\rVert$, and be left therefore with optimizing only with respect to the direction $\bm{\hat{a}}\coloneqq\bm{a}/\left\lVert\bm{a}\right\rVert$. In formulas
@@
Equation (13) is a first order nonlinear PDE belonging to the well known class of static Hamilton-Jacobi PDEs [2, 27, 1]. In general, it does not admit a closed-form solution, however, very efficient algorithms exist to solve it approximately. Before moving on let us briefly comment on $g$. In our derivation we supposed that the aircraft flies at the speed that yields the optimal glide ratio, implying that the pilot should adapt the magnitude of the airspeed to the local wind. Alternatively, to obtain more conservative estimates one can assume instead that the aircraft flies at a constant horizontal airspeed $v$. In this case $g(\bm{x},z,\bm{\hat{a}})=\left\lVert\tilde{\bm{a}}+\bm{W}(\bm{x},z)\right\rVert/s(v)$, where $\tilde{\bm{a}}$ solves $\bm{\hat{a}}=\frac{\bm{W}(\bm{x},z)+\tilde{\bm{a}}}{\left\lVert\bm{W}(\bm{x},z)+\tilde{\bm{a}}\right\rVert}$, with $\left\lVert\tilde{\bm{a}}\right\rVert=v$.
@@ II.3 Gliding as front propagation
which is known as the eikonal equation. The initial condition is $U(\bm{x}_{0})=z_{0}$. The eikonal equation arises when describing the arrival times of a front at each point in space. Suppose we have a sheet of paper and at time $t=0$ we set fire to it at a point $\bm{x}_{0}\in\operatorname{\mathbb{R}}^{2}$. Suppose further that the fire front advances through the paper with speed $g$. Then letting $U(\bm{y})$ be the time at which the fire front reaches $\bm{y}$, we have that $U$ satisfies (16). We can also include obstacles in this analogy. Suppose some portion of the paper is wet. In this case, the fire front stops at the border of the wet region. This is similar to what happens when the glider’s altitude falls below $h_{\text{min}}$. This analogy shows how we can see the descending glider as an expanding front that passes around the obstacles until it collides with them and stops.